OpenStudy (anonymous):

Medal and follow .. Choose the equation of the line passing through the point (-4, -2) and perpendicular to y=-x + 6

OpenStudy (anonymous):

do you need it in slope-intercept form or point-slope form?

OpenStudy (anonymous):

Ill show answers

OpenStudy (anonymous):

A y =-x + 6 B y= C y= D y=

OpenStudy (anonymous):

B y=-x-6

OpenStudy (anonymous):

C = -x - 2

OpenStudy (anonymous):

D = x+ 2

OpenStudy (anonymous):

In b there is no -in front of x

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Okay so the slope of the perpendicular line will be -x because the reciprocal of -1 is -1. so now since we do not have the y-intercept we cant put it into that form yet so we use point-slope form: y - y1 = m(x - x1) plug in the points: y-(-2) = -1(x-(-4)) Simplify: y+2=-1(x+4) Distribute the -1: y+2=-x-4 We have to get y alone so subtract 2 from each side: y= -x-6 s

OpenStudy (anonymous):

Is it c

OpenStudy (anonymous):

can you repost what answer a is?

OpenStudy (anonymous):

Y= -x + 6

OpenStudy (anonymous):

Hello?!?!

OpenStudy (anonymous):

@ChihuahuaAdorer

OpenStudy (anonymous):

I think it is c

OpenStudy (anonymous):

Oh ok c is -x - 2

OpenStudy (anonymous):

yes